# HyperMOOC: Augmenting MOOC Videos with Concept-based Embedded Visualizations

### v1

@@ 5.2. Expert Interviews
6. Future Work and Limitations
Significance. Currently, communicating data through video-based visualizations remains a challenging task (Thompson et al., 2020), where online learning videos, as the most popular means of distance learning, lack effective methods for learners to quickly and effectively access data directly from videos. Our work provides a way for learners to seamlessly access the data they are interested in. On the other hand, numerous studies on online learning experiences highlight the importance of interface design, learning design, and the online context in shaping students’ engagement and learning behaviors (Wong et al., 2024; Wong and Hughes, 2023). By employing user-friendly visual interfaces to stimulate learner interest and motivation, we offer a potential direction for improving the effectiveness and experience of online learning.
Accessibility. In recent years, eye-tracking technology has become increasingly affordable. Compared to other interaction mediums such as keyboard, mouse, and voice, gaze input allows for rapid and intuitive interactions. Many gaze interaction methods are used in visual analytics tasks (Silva et al., 2019). However, our tool only supports using mouse hover to replace the most basic Gaze Focus. In the future, more advanced gaze interactions (e.g., fixation, saccade, pursuit) should be incorporated (Chen et al., 2023; Silva et al., 2018). Another promising direction is combining gaze interactions with other interaction mediums (such as voice and gesture) to achieve more personalized embedded visualizations. Moreover, since MOOC platforms are typically online, our tool should be easy to integrate, i.e., learners can ignore configuration, data pipelines, and platforms to use our tool. Plug-ins (Nobarany, 2014; Wolfinger, 2008) is a good means to achieve this. Therefore, HyperMOOC should offer a plugin version to support rapid integration into various MOOC platforms.
Generalizability. Although our prototype tool focuses on augmenting MOOC videos, once the algorithms and datasets become available, its design can be extended to other educational materials, such as debates (South et al., 2020) and presentations (Zeng et al., 2019). Furthermore, our design practices can be applied within courses to augment discussions, interactions, or even potentially indicate good study habits among learners, further promoting an immersive classroom experience. Moreover, as augmented reality (AR) and virtual reality (VR) are emerging as mainstream platforms (Squires, 2017),(Cortiz and Silva, 2017) for future learning, extending our approach to AR/VR-based learning environments presents a promising direction for exploring how embedded visualizations can enhance the learning experience in immersive settings.
Study limitations. As the first attempt to augment MOOC videos, HyperMOOC, similar to other research on augmented videos (Chen et al., 2023), has several limitations. Firstly, our evaluation involved a small sample size of participants and a limited number of domain experts. While we have made progress in assessing immediate learner outcomes through post-video quizzes, we did not evaluate long-term knowledge retention and transfer. Secondly, our design space was derived from a limited number of existing literature reviews, primarily identifying common behavior patterns among MOOC video learners. The current design space only includes learning content and learners, without considering the potential impact of video creators’ (instructors’) teaching habits on learning performance. Lastly, we recognize the need to consider more augmentable elements to promote classroom engagement. Future research should expand the sample size, incorporate long-term learning effect assessments, and explore a more comprehensive design space, including the influence of teaching methods on learning. These improvements will help further optimize MOOC video augmentation systems and enhance their effectiveness across various learning contexts.

### v2

@@ 5.2. Expert Interviews
6. Discussion
We provide a comprehensive analysis of HyperMOOC’s modes, discuss the significance and generalization of our findings, and outline limitations and future directions.
6.1. Comprehensive Analysis
The study evaluates HyperMOOC using three quantitative measures: test scores, SUS usability, and NASA-TLX workload. Overall, the FULL mode outperforms both RAW and AUG across all metrics—learners achieved higher test scores, rated the system as more usable, and mostly reported lower cognitive load. This advantage is particularly evident for learners unfamiliar with the content, suggesting that additional structure and guidance translate into substantial learning gains for these groups. However, no single mode is universally optimal: a subset of learners preferred AUG for its simpler interface while preserving key annotations, indicating that excessive embedded information may impose additional cognitive burden for complete beginners.
From the perspective of Cognitive Load Theory (CLT) (Sweller et al., 1998, 2019; Mutlu-Bayraktar et al., 2019), traditional video designs with external panels require learners to switch between interfaces while mentally reconstructing content, increasing extraneous load (Jeuris and Bardram, 2016). FULL mode addresses this by embedding supporting information directly into the video frame—key concepts, highlights, and navigation links are placed near explanations, reducing task-irrelevant operations and helping learners allocate more effort to comprehension (Alpizar et al., 2020; Xie et al., 2017). The three-stage interaction flow (Play, Focused, Paused) makes an intuitive learning rhythm explicit, mitigating overload from presenting too much information at once (Spanjers et al., 2012; Liu et al., 2024). From a pedagogical perspective, embedding element highlights and concept-based segmentation cues acts as instructional scaffolding to transform passive viewing into active processing (Alpizar et al., 2020; Xie et al., 2017; Liu et al., 2024). Adhering to the spatial contiguity principle, in-situ navigation minimizes split-attention effects and preserves conceptual continuity (Mutlu-Bayraktar et al., 2019). This facilitates metacognitive monitoring by enabling seamless, self-directed watch–review–search cycles, redirecting cognitive resources toward constructing deeper mental models rather than mere retrieval (Sweller et al., 1998, 2019; Jeuris and Bardram, 2016).
Expert interviews complement these quantitative findings. Both experts emphasized three strengths of HyperMOOC: (1) clearer structural cues for concept relationships; (2) key prompts and navigation embedded within the video, reducing interface switching; and (3) multi-stage interaction enabling on-demand exploration rather than passive reception. They also noted a boundary: for learners with weaker backgrounds, the information density in FULL can feel overwhelming initially, suggesting that a more gradual introduction or simpler default view may be needed—consistent with the preference variability in our data.
6.2. Significance & Generalization
HyperMOOC applies principles from cognitive load theory, signaling, and segmented learning (Sweller et al., 1998, 2019; Mutlu-Bayraktar et al., 2019; Alpizar et al., 2020; Spanjers et al., 2012; Liu et al., 2024) to hypervideo interface design. By embedding concept prompts and navigation links directly into the video frame, HyperMOOC reduces the cognitive cost of interface switching (Jeuris and Bardram, 2016), offering a concrete example of translating learning theory into interactive design patterns. The video is thus designed as a cognitive interface that supports and guides learning, rather than serving merely as a passive delivery medium. From an HCI perspective, HyperMOOC demonstrates how embedded hypervideo can reallocate attention in information-dense learning tasks, with video serving as an active participant in organizing users’ cognitive activities.
The design method behind HyperMOOC shows how to decompose design problems along multiple dimensions (data types, visual representations, user behaviors, cognitive states) and recombine them into concrete interaction strategies, then instantiate and compare different variants (RAW, AUG, FULL) (Sauli et al., 2018; Ortiz and Moya, 2024). This “decompose–recombine–evaluate” approach generalizes beyond MOOC videos to other information-intensive systems such as programming tutorials, professional training, data-driven storytelling, and AR/VR learning environments.
Based on these results, we summarize three design insights applicable to broader human-computer interaction and educational contexts beyond this system:
  • Progressive Cognitive Scaffolding. Information hiding goes beyond UI cleanup; it is a pedagogical strategy to manage intrinsic load. By presenting essential concepts first and details only on demand, systems create a manageable learning zone. This structure prevents overload and encourages active inquiry, acting as a scaffold to support learners’ gradual independence (Spanjers et al., 2012; Liu et al., 2024).
  • Adaptive Knowledge Layering. Systems should provide varying ”thickness” of information to match learner expertise. Novices need explicit guidance for schema construction, while experts benefit from concise summaries. Providing flexible information levels mitigates the expertise reversal effect, ensuring that instructional support remains effective rather than redundant as learners progress (Sweller et al., 1998, 2019; Mutlu-Bayraktar et al., 2019).
  • Situated Active Learning. Placing navigation directly within the video adheres to the spatial contiguity principle, effectively minimizing split-attention. This turns passive video viewing into an active workspace. Such interaction strengthens dual coding, helping learners integrate visual and auditory information into a unified mental model without the cost of context switching (Ginns, 2006; Jeuris and Bardram, 2016; Ortiz and Moya, 2024).
These insights are mainly derived from courses with relatively dense concepts and clear structure. In more open-ended, discussion-driven, or inspiration-oriented content, lighter-weight augmentation may be more suitable(Sauli et al., 2018). More broadly, these strategies generalize to other information-rich learning environments. In immersive learning (AR/VR), the in-situ principle mitigates spatial disorientation, while Intelligent Tutoring Systems (ITS) can leverage adaptive layering to automate support fading. They also motivate further work in other domains to examine how information hierarchies should be defined, how display strategies can be adapted dynamically to user characteristics, and how the idea of contextual navigation can be extended to non-video complex interfaces. Ultimately, framing interface elements as cognitive supports offers a blueprint for cognitive interfaces that bridge technical scalability with pedagogical depth.
6.3. Future Work & Limitations
Assessment. Our work is a first attempt to augment MOOC videos with concept-based embedded visualizations, yet presents limitations. First, regarding our sample ($N=36$ learners, $N=4$ experts), while it aligns with contextual practices in related work (Chen et al., 2023) and provided sufficient statistical power, the size limits broad generalizability. Future studies should recruit larger, diverse cohorts. Second, using academic major as a proxy for prior knowledge may confound knowledge level with quantitative skills or visualization literacy. Third, our metrics focus on immediate outcomes and subjective experience, overlooking long-term retention or transfer. Finally, the design primarily reflects learner behaviors; future work should include instructor-focused studies to better connect pedagogical intent with system design.
Interaction modalities. Our current prototype uses mouse hover to approximate basic Gaze Focus behavior. While prior literature suggests mouse movement is a strong correlate of visual attention (Huang et al., 2011), we acknowledge that it is not a perfect substitute for precise gaze data. The lack of eye-tracking verification in the current study is a limitation, as mouse movements may occasionally decouple from actual visual focus. With eye-tracking technologies becoming increasingly affordable (Silva et al., 2019), future work should incorporate richer gaze-based techniques—such as fixation, saccade, and smooth pursuit (Chen et al., 2023; Silva et al., 2018)—to better support learner interaction with embedded visualizations. Another promising direction is combining gaze with other modalities (e.g., voice and gesture) to provide more inclusive control and investigate how different modality combinations influence usability and learning outcomes.
Pipeline generalizability. The data processing pipeline (Fig. 5) adopts a modular, bottom-up architecture to facilitate generalization. Most components are domain-agnostic: SBD/EMD for slide detection, Google Speech Recognition for audio transcription, OpenFace and TextRank for auxiliary element extraction, and TOT for course structure modeling. At the event level, GPT-3.5 with customizable prompts enables adaptation to different subject domains without model retraining. However, certain limitations exist. The YOLOX model for element detection was trained on frames across multiple disciplines, but specialized visual content (e.g., medical imaging, musical notation) may require domain-specific fine-tuning. Additionally, our pipeline is primarily validated on slide-based lecture formats; non-slide formats such as laboratory demonstrations, field recordings, or whiteboard-only lectures may require extensions for element detection. Furthermore, concept-intensive STEM courses align well with our hierarchical data representation, whereas discussion-based humanities courses with less explicit concept structures may require alternative relationship modeling approaches. Future work should evaluate pipeline performance across broader MOOC platforms, content formats, and language contexts to establish generalization boundaries.
Adaptive visualization. Our results show substantial variation in how learners respond to different augmentation levels in terms of information density and perceived complexity, highlighting adaptive visualization as a key future direction. This involves determining how much information and which interaction capabilities should be exposed at a given moment—including concept density, relationship richness, visual encoding complexity, and advanced interactions such as filtering and path tracing. Future work can formalize complexity tiers mapping these dimensions to clear visualization configurations, and investigate tier transitions within videos and across courses. Adaptation should also account for prior knowledge differences: novices benefit from thinner information layers and stronger guidance, while advanced learners may prefer richer representations. Future systems may combine system-driven adaptation (adjusting based on pre-test scores or performance) with learner-driven customization (choosing between basic and advanced views). Furthermore, incorporating direct cognitive load indicators and physiological signals into the adaptive loop could enable real-time complexity adjustments—reducing visible elements when load is high and offering richer visualizations when load is low. Designing and validating such cognitive-load-aware adaptive strategies constitutes a substantial research agenda spanning visualization, learning analytics, and HCI.
@@ 7. Conclusion
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China.(Nos. 62277013 and 62177040), the National Social Science Fund Major Project (No. 24&ZD075), the Zhejiang Lingyan Project (No. 2026C02A1238), the Central Government-Guided Local Science and Technology Development Fund (Grant No. 2025ZY01045) and the Zhejiang Provincial Key Laboratory of Brain Computer Collaborative Intelligence Technology and Applications(Grant No.2025E10015).